# Tiny-BioMoE: a Lightweight Embedding Model for Biosignal Analysis

### v5

@@ 2. Related Work
In addition, multimodal approaches combining behavioral and physiological data have gained increasing attention in recent years, with several studies demonstrating the benefits of integrating multiple sources of information to improve performance. Many researchers explore combination of biosignals due to their potential to provide more accurate measurements. For instance, Jiang et al. (Jiang et al., 2024a) proposed a hybrid temporal-channel attention model that fuses ECG and galvanic skin response (GSR) signals, while in (Jiang et al., 2024b), the authors employed ECG and GSR signals alongside neural networks enhanced with Squeeze-and-Excitation blocks. Chu et al. (Chu et al., 2017) integrated blood volume pulse, ECG, and skin conductance level, using a hybrid approach that combined genetic algorithm-based feature selection with principal component analysis. On the other hand, Badura et al. (Badura et al., 2021) employed a multimodal setup in their effort to assess pain during physiotherapy, which included EDA, EMG, respiration, blood volume pulse (BVP), and hand force. Other researchers focus on combining physiological with behavioral modalities. The authors in (Zhi and Yu, 2019) utilized facial videos in combination with ECG, EMG, and EDA, extracting a wide range of handcrafted features and fusing them. In (Gkikas et al., 2024), the authors combined facial videos and heart rate data to develop a transformer-based framework for recognizing pain intensity. Similarly, in (Farmani et al., 2025), CNN and LSTM models were used to analyze facial videos and EDA, achieving high performance. Finally, the authors in (Gkikas et al., 2025c) proposed a foundation model that extracted feature representations from a wide range of modalities—both behavioral and physiological—for pain assessment.
@@ 4. Experimental Evaluation & Results
This study leverages the dataset released by the challenge organizers, which consists of electrodermal activity recordings from $65$ participants. Data collection took place at the Human-Machine Interface Laboratory, University of Canberra, Australia, and is divided into $41$ training, $12$ validation, and $12$ testing subjects. Pain stimulation was induced using transcutaneous electrical nerve stimulation (TENS) electrodes positioned on the inner forearm and the back of the right hand. Two pain levels were measured: pain threshold—the minimum stimulus intensity perceived as painful (low pain), and pain tolerance—the maximum intensity tolerated before becoming unbearable (high pain). The signals have a frequency of $100$ Hz and a duration of approximately $10$ seconds. We refer to (Fernandez Rojas et al., 2023b) for a detailed description of the recording protocol and to (Fernandez Rojas et al., 2024a) for information regarding the previous edition of the challenge. This study utilises all available modalities—EDA, BVP, respiratory signals, and peripheral oxygen saturation (SpO2). All experiments were conducted on the validation subset of the dataset and evaluated within a multi-class classification framework, encompassing three pain levels: No Pain, Low Pain, and High Pain. The validation results are reported in terms of macro-averaged accuracy, precision, and F1 score. The final results of the testing set are also reported. We note that all experiments followed a deterministic setup, eliminating the effect of random initializations; thus, any performance differences arose strictly from the chosen optimization settings, modalities, or other intentional changes rather than chance.
@@ 6. Conclusion
This work relied on the AI4PAIN dataset (Fernandez Rojas et al., 2023b, 2024a), made available by the challenge organizers, to assess automatic pain recognition methods. All participants confirmed the absence of neurological or psychiatric conditions, unstable health issues, chronic pain, or regular medication use during the session. Before the experiment, participants were thoroughly informed of the procedures, and written consent was obtained. The original study’s human-subject protocol received ethical clearance from the University of Canberra’s Human Ethics Committee (approval number: 11837). The proposed method was developed for continuous pain monitoring, aiming to enhance pain assessment protocols and improve patient care. However, as validation and testing were performed on controlled laboratory data, its deployment in real-world clinical settings requires further investigation and comprehensive evaluation.

### v6

@@ 2. Related Work
In addition, multimodal approaches combining behavioral and physiological data have gained increasing attention in recent years, with several studies demonstrating the benefits of integrating multiple sources of information to improve performance (Farmani et al., 2025a). Many researchers explore combination of biosignals due to their potential to provide more accurate measurements. For instance, Jiang et al. (Jiang et al., 2024a) proposed a hybrid temporal-channel attention model that fuses ECG and galvanic skin response (GSR) signals, while in (Jiang et al., 2024b), the authors employed ECG and GSR signals alongside neural networks enhanced with Squeeze-and-Excitation blocks. Chu et al. (Chu et al., 2017) integrated blood volume pulse, ECG, and skin conductance level, using a hybrid approach that combined genetic algorithm-based feature selection with principal component analysis. On the other hand, Badura et al. (Badura et al., 2021) employed a multimodal setup in their effort to assess pain during physiotherapy, which included EDA, EMG, respiration, blood volume pulse (BVP), and hand force. Other researchers focus on combining physiological with behavioral modalities. The authors in (Zhi and Yu, 2019) utilized facial videos in combination with ECG, EMG, and EDA, extracting a wide range of handcrafted features and fusing them. In (Gkikas et al., 2024), the authors combined facial videos and heart rate data to develop a transformer-based framework for recognizing pain intensity. Similarly, in (Farmani et al., 2025b), CNN and LSTM models were used to analyze facial videos and EDA, achieving high performance. Finally, the authors in (Gkikas et al., 2025c) proposed a foundation model that extracted feature representations from a wide range of modalities—both behavioral and physiological—for pain assessment.
@@ 4. Experimental Evaluation & Results
This study leverages the dataset released by the challenge organizers, which consists of electrodermal activity recordings from $65$ participants. Data collection took place at the Human-Machine Interface Laboratory, University of Canberra, Australia, and is divided into $41$ training, $12$ validation, and $12$ testing subjects. Pain stimulation was induced using transcutaneous electrical nerve stimulation (TENS) electrodes positioned on the inner forearm and the back of the right hand. Two pain levels were measured: pain threshold—the minimum stimulus intensity perceived as painful (low pain), and pain tolerance—the maximum intensity tolerated before becoming unbearable (high pain). The signals have a frequency of $100$ Hz and a duration of approximately $10$ seconds. We refer to (Fernandez Rojas et al., 2025, 2023b) for a detailed description of the recording protocol and to (Fernandez Rojas et al., 2024a) for information regarding the previous edition of the challenge. This study utilises all available modalities—EDA, BVP, respiratory signals, and peripheral oxygen saturation (SpO2). All experiments were conducted on the validation subset of the dataset and evaluated within a multi-class classification framework, encompassing three pain levels: No Pain, Low Pain, and High Pain. The validation results are reported in terms of macro-averaged accuracy, precision, and F1 score. The final results of the testing set are also reported. We note that all experiments followed a deterministic setup, eliminating the effect of random initializations; thus, any performance differences arose strictly from the chosen optimization settings, modalities, or other intentional changes rather than chance.
@@ 6. Conclusion
This work relied on the AI4PAIN dataset (Fernandez Rojas et al., 2023b, 2024a, 2025), made available by the challenge organizers, to assess automatic pain recognition methods. All participants confirmed the absence of neurological or psychiatric conditions, unstable health issues, chronic pain, or regular medication use during the session. Before the experiment, participants were thoroughly informed of the procedures, and written consent was obtained. The original study’s human-subject protocol received ethical clearance from the University of Canberra’s Human Ethics Committee (approval number: 11837). The proposed method was developed for continuous pain monitoring, aiming to enhance pain assessment protocols and improve patient care. However, as validation and testing were performed on controlled laboratory data, its deployment in real-world clinical settings requires further investigation and comprehensive evaluation.